Find the polynomial for the perimeter and area.The Width is b+3 and the length is b.

Answer by checkley77(12844)

2L+2W=PERIMETER
2b+2(b+3)=perimeter
2b+2b+6=perimeter
4b+6=perimeter.
------------------
LW=area
b(b+3)=area
b^2+3b=area.
------------------